OpenStudy (anonymous):

write in point slope form the equation of the line that passes through the given point and has the given slope (-1, 0) m= 3 I GIVE MEDALS

OpenStudy (anonymous):

D you know the formula for point slope?

OpenStudy (anonymous):

y-y1 = m (x-x1)

OpenStudy (anonymous):

0-y1=31(-1-x1)????